Haemoglobin is a chemical in blood that is responsible for transporting oxygen from the lungs to other parts of the body a solution containig 1g of Haemoglobin in 1000cm³ of water had an osmotic preasure of 365N/m² at 298k estimate the molar mass of the Heamoglobin
1
Expert's answer
2019-06-24T04:21:20-0400
"\pi = MRT"

"\pi = 365 Pa = 365 Pa\times \frac{1 atm}{101325 Pa} =0.003602 atm"


"M = \frac{n}{V} = \frac{moles}{1 L}"


"R = 0.0821\frac{L\times atm}{mol \times K}"


"T = 298 K"


"0.003602 atm = \frac{n}{1L} \times 0.0821 \frac{L\times atm}{mol \times K} \times 298 K"

"n = 1.47\times 10^{-4} moles"

"m = M\times n"


"\therefore M = \frac{m}{n} = \frac{1 g}{1.47\times 10^{-4}} = 6792 \frac{g}{mol} =6790 \frac{g}{mol}"

Answer: "6790 \frac{g}{mol}"